OpenStudy (anonymous):

The area of a rectangle is given by A = x2 – 2x + 30. Find the value of x if the total area of the rectangle is equal to 45 square units.

OpenStudy (anonymous):

x^2 - 2x + 30 = 45 x^2 - 2x - 15 = 0 Factoring gives: (x+3)(x-5) = 0 Solutions: x = -3 OR x = 5. Since a side of rectangle can't be negative, take 5 as the value of x.

OpenStudy (anonymous):

It says find value of x. So, x value is 5.